Listed below are paired data consisting of movie budget amounts and the amounts that the movies grossed. Find the regression​ equation, letting the budget be the predictor​ (x) variable. Find the best predicted amount that a movie will gross if its budget is ​$140 million. Use a significance level of α=0.05.

Budget ($) in Millions   Gross ($) in Millions
46                                 107
19      16
115   94
72                                   71
73      120
54                                  109
119      95
62                                  105
8      57
61   112
122      203
17                                     43
13      12
147   295
10                                      53

The regression equation is ^y = ? + ?x
(Round to one decimal place as​ needed.)

The best predicted gross for a movie with a ​$140 million budget is ​$? million.
​(Round to one decimal place as​ needed.)
Answer #1

using minitab>stat>Regression

we hav e

Regression Analysis: gross versus budget

The regression equation is
gross = 20.23 + 1.267 budget


S = 44.0907 R-Sq = 65.1% R-Sq(adj) = 62.4%


Analysis of Variance

Source DF SS MS F P
Regression 1 47065.9 47065.9 24.21 0.000
Error 13 25271.9 1944.0
Total 14 72337.7

The regression equation is y =20.2 + 1.3*x

The best predicted gross for a movie with a ​$140 million budget is
y =20.2 + 1.3*140 = 202.2